OpenStudy (anonymous):

please help me ... divide:(y^2-y-12)/(y^2-6y+8 by (y^2+2y-15)/(y^2-7y+10)

OpenStudy (anonymous):

\[\frac{y^2-y-12}{y^2-6y+8}\times \frac{y^2-7y+10}{y^2-2y-15}\] is a start then factor then cancel

OpenStudy (anonymous):

ok thanks so much

OpenStudy (anonymous):

yw good luck with the factoring

OpenStudy (anonymous):

i am good at canceling but i need help with factoring can u plz help me with the factoring

OpenStudy (anonymous):

@dumbcow can u please help me factor this

OpenStudy (anonymous):

@Compassionate

OpenStudy (dumbcow):

"i am good at cancelling" haha sorry but thats funny

OpenStudy (anonymous):

it is ok i am stupid when it comes to math

OpenStudy (compassionate):

Factoring just means breaking it down into smaller parts For example, lets take the first one. \[x^2 - y - 12\] (x )(x ) Because x * x = x^2 Then we wil lsay, "what two products multiplied together equal 12, and added equal -y? Well, you'll find it's not factorable because no two terms can be multiplied but their product be -y.

OpenStudy (dumbcow):

so for factoring you want to find factors of constant which add up to middle coefficient \[y^{2} + by + c = (y+e)(y+f)\] e*f = c e+f = b

OpenStudy (anonymous):

oh ok i see